OpenStudy (anonymous):

Simplify the radical expression...... 3x^2 - 7x + 4 / 5x^2 - 8x + 3

OpenStudy (anonymous):

\[\frac{4-3 x}{3-5 x} \]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

factor top and bottom and cancel

OpenStudy (anonymous):

?????????????

OpenStudy (anonymous):

I'm sorry, I have to get up for work in 21/2 hours and my brain is completely fried!

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

numerator factors to (x-1)(3x-4) denom factors to (x-1)( 5x-3) x-1 cancels, leaving (3x-4) / (5x-3)

OpenStudy (anonymous):

\[\frac{((-1+x) (-4+3 x))}{((-1+x) (-3+5 x))} \]or\[\frac{3 x-4}{5 x-3} \]

OpenStudy (anonymous):

tytytyty both soooo much...............